Suppose a normally distributed numerical variable X has MU = 15 and Sigma = 6. Answer the following questions about the sampling distribution of the mean if the sample size is 100.

1. The sampling distribution of X bar is (blank) distributed with mu X bar = (blank) and sigma X bar = (blank). (fill in the blanks)

2. Suppose a random sample is chosen. what is the probability that this selected sample mean is less than 14.2?

3. What is the probability that the selected sample mean is between 15 and 16?
Answer #1

solutions - el = 15 in I Noormally I distributed lize= 115 2 = x - 2) p (fc14.2) = Plac 14.2-15 0.6 =plz (-1.43) - 10.0764 |